# Optimal ambition in business, politics and life

### v2

@@ II.3 Social comparison hurts performance and penalizes ambition
We have previously assumed that agents know the reward distribution. What if they instead estimate the distribution and evaluate their satisfaction by looking at peers’ rewards Goethals and Darley (1977)—social comparison Suls et al. (2002); Schulz (2015); Lévy-Garboua and Montmarquette (2004); Bygren (2004)? Indeed, people often focus specifically on those performing better—‘upward social comparison’ Gerber et al. (2018); Muller and Fayant (2010).
We model cohort comparison and upward social comparison. Cohort comparison assumes that each agent knows the mean and variance of their peers’ rewards at the previous time step. Upward social comparison assumes that each agent makes these estimates only taking into account their own reward and peers’ rewards that were greater. Note that when agents define their satisfaction threshold relative to the performance of others, a previously satisfied agent can become unsatisfied and return to searching if group performance changes compared to their threshold.
Fig. 4 shows simulations for 100-agent cohorts. Agents vary in their satisfaction thresholds, each chosen randomly from a uniform distribution. The optimal satisfaction threshold is slightly lower with cohort comparison than when agents know the true distribution. Because agents can become unsatisfied if the cohort statistics change, setting a lower threshold prevents agents from overreacting to others’ success.
@@
Figure 4: Social comparison is costly. Mean rewards for cohort comparisons and upward social comparisons are shown on a rough, non-skewed landscape. Cohort comparison (yellow curve) lowers the optimal threshold and cumulative reward. Upward social comparison (purple curve) further lowers the optimal threshold and hinders performance. Both expected reward and optimal satisfaction threshold are lower compared to the reference landscape (black, dashed curve). $\epsilon_{t}\sim\mathcal{N}(\mu,\sigma^{2})$, $\varphi=0$, $t_{\text{max}}=1000$, cohorts of $100$ agents, with thresholds uniformly sampled from $[-3,6]$, averaged over $10^{4}$ simulations.
@@ II.4 Applications
Fig. 5 illustrates examples where our model’s insights could be applied. Economic growth tends to be left-skewed in developed countries Burgess et al. (2020b) (Fig. 5B shows U.S. data.) Private wealth distributions are right-skewed Scheffer et al. (2017) (Fig. 5B). As a result, risk taking is often beneficial in business and private investment Benuzzi and Ploner (2024), but costly in policymaking. However, our model implies seemingly opposite advice for ambition, compared to the mean. Mean business success is highly influenced by rare and extremely successful ‘unicorns’. Thus, an entrepreneur who sets a high satisfaction thresholds compared to the mean may miss opportunities for important and achievable successes. A country’s mean economic growth rate is dragged down by large and rare recessions, so policymakers would be unwise to be satisfied with growth rates close to this mean in a typical year.
@@
Fig. 5C shows the skewness and smoothness of monthly polling margins in swing states ahead of the U.S. 2020 presidential election Five Thirty Eight (2020), from the perspective of the Republican party (for Democrats, skewness would have the opposite sign and smoothness would be the same). Higher margins are more likely in states with more rugged and left-skewed polling distributions. Our model implies that parties should pursue more ambitious (albeit less risky, given the left-skew) political strategies in such states.
@@
Figure 5: A: U.S. growth rates in gross domestic product (GDP) per capita (gray) Our World in Data (2024), compared moving means (black) and medians (dashed). The median is higher than the mean, indicating left skew. B: distributions of private wealth among the 2500 top companies Dhiman (2024) and billionaires Elgiriyewithana (2023). The means are much greater than the medians, indicating right skew. C: monthly polling margins, from the perspective of Republicans (i.e. Republican share - Democrat share) in U.S. swing states during 2020 prior to the presidential election Five Thirty Eight (2020). Darker colors indicate bigger average differences between the two parties. Lighter colored states have smaller average margins.
@@ III Discussion
Our model formalizes the folk intuition that optimal ambition is both finite and aims to do better than average (Fig. 2A). We show empirical examples (dating and college applications) in which people seem to behave as though they understand this (Fig. 2B,C). Although our model assumes all agents face the same reward distribution, the available rewards—and consequently, the optimal ambition—varies widely across people and contexts in reality.
@@
Our model then adds more precise insights. First, we show being too unsatisfiable is costlier than being too easily satisfied, on average (Fig. 2A). Second, we show that left-skewed and rugged reward distributions should motivate more ambition (Fig. 3). This exposes a subtle difference between ambition and risk, as left-skewed rewards motivate less risk-taking Benuzzi and Ploner (2024). Third, we show that upward social comparison is costly by creating unrealistic expectations (Fig. 4). We describe how these insights apply to real-world contexts relating to politics, economic policy, and wealth building (Fig. 5), The introduction discussed additional examples.
@@
Our model is admittedly simple, though ‘explore-exploit’ and sequential searches are common abstractions in areas such as harvesting Richard et al. (2018), animal foraging MacArthur and Pianka (1966); Cowie (1977), and management science Lavie et al. (2010); Greve (2007). Our ‘satisfaction’ framework is somewhat related to the satisficing concept in behavioral economics Simon (1955), but it is also distinct in that our threshold is optimized.
@@
Nonetheless, future research should explore complexities our model does not consider. For example, there can be inherent benefits to sticking with one strategy (e.g., accruing expertise, deepening a relationship) or costs to moving on or exploring (e.g., disruptions from moving cities or changing jobs). Real people are also prone to weighing potential losses more than potential gains (‘loss aversion’ Kahneman et al. (1991)). This would reduce the optimal satisfaction threshold in absolute terms, by reducing the utility of low-end rewards. Our agents are simplified both in terms of their decisions and their inferences about available rewards. Our social comparison model does not consider structured social networks. We model random rather than intelligent search processes, though it is straightforward to conjecture that intelligent search would increase optimal ambition by making high, but rare, rewards easier to find.
@@
Despite these limitations, our conceptual results are intuitive and likely to generalize. For example, under any search process, increasing ambition makes satisfactory rewards harder to find. Making high rewards more abundant—via skewness or ruggedness—makes ambitious targets easier to achieve. Upward social comparison distorts perceptions of reality, which hinders rational decision making.
@@
How ambitious to be is important in a wide range of professional, political, and personal contexts. Our model, results, and examples offer entry points for mathematically precise but intuitive explorations of this question.
IV Materials and Methods

### v3

@@ II.3 Social comparison hurts performance and penalizes ambition
We have previously assumed that agents know the reward distribution. What if they instead estimate the distribution, and evaluate their satisfaction, by looking at peers’ rewards Goethals and Darley (1977)? This type of behavior is called social comparison Suls et al. (2002); Schulz (2015); Lévy-Garboua and Montmarquette (2004); Bygren (2004). Indeed, in reality people often focus specifically on peers receiving higher rewards than they are—‘upward social comparison’ Gerber et al. (2018); Muller and Fayant (2010).
We model cohort comparison and upward social comparison. Cohort comparison assumes that each agent knows the mean and variance of their peers’ rewards (across all peers) at the previous time step. Upward social comparison assumes that each agent makes these estimates only taking into account their own reward and peers’ rewards that were greater. Note that when agents define their satisfaction threshold relative to the performance of others, a previously satisfied agent can become unsatisfied and return to searching if group performance changes compared to their threshold.
Fig. 4 shows simulations for 100-agent cohorts. Agents vary in their satisfaction thresholds, each chosen randomly from a uniform distribution. The optimal satisfaction threshold is slightly lower with cohort comparison than when agents know the true reward distribution. Because agents can become unsatisfied if the cohort statistics change, setting a lower threshold prevents agents from overreacting to others’ success. Additional cohort sizes and threshold intervals are shown in Fig. A1 and Fig. A2.
@@
Figure 4: Social comparison is costly. Mean rewards for cohort comparisons and upward social comparisons are shown on a rough, non-skewed landscape. Cohort comparison (yellow curve) lowers the optimal satisfaction threshold and cumulative reward. Upward social comparison (purple curve) further lowers the optimal threshold and substantially hinders performance. Both the expected reward and the optimal satisfaction threshold are lower compared to the reference landscape (black, dashed curve), which assumes no social comparison. Parameter values are: $\epsilon_{t}\sim\mathcal{N}(\mu,\sigma^{2})$, $\varphi=0$, and $t_{\text{max}}=1000$. Cohorts include $100$ agents, with thresholds uniformly sampled from $[-3,6]$. Results shown are averaged over $10^{4}$ simulations.
@@ II.4 Applications
Fig. 5 illustrates examples of where our model’s insights regarding smoothness and skewness could be applied to economic policy, entrepreneurship, and political campaigns. These examples reveal a subtle but important distinction between optimal ambition and optimal risk-taking.
Economic growth rates tend to have left-skewed distributions in developed countries, because recessions tend to have larger magnitudes than booms Burgess et al. (2020b). Fig. 5a shows this pattern in U.S. data. From a policy standpoint, the left-skewed distribution of economic growth rates implies that risk taking is costly. If downside risks are larger than upside risks, risk-taking lowers the expected outcome. However, our model implies seemingly opposite advice for ambition, compared to the mean reward. A country’s mean economic growth rate is dragged down by large and rare recessions (Fig. 5a). Policymakers would therefore be unwise to be satisfied with growth rates close to this mean in a typical year. In other words, the left-skewness of economic growth rates calls for less risky but more ambitious (compared to the mean) economic policymaking.
Private wealth distributions have the opposite property: they are heavily right-skewed Scheffer et al. (2017). Fig. 5b illustrates the wealth distributions of the world’s richest billionaires and the world’s largest companies. A right-skewed distribution of private wealth outcomes implies that upside risks are often larger than downside risks—risk-taking in entrepreneurship often means either winning big or losing small. Consequently, risk taking is often beneficial in business and private investment Benuzzi and Ploner (2024). However, our model suggests that the right skewness of private wealth calls for less ambition, compared to the mean available reward. This is counterintuitive, but is related to the fact that mean success in wealth building and entrepreneurship is highly influenced by rare and extremely successful ‘unicorns’. Thus, an entrepreneur or investor who sets a high satisfaction threshold compared to the mean may miss opportunities for important and achievable successes.
@@
Fig. 5c shows the skewness and smoothness of monthly polling margins in swing states ahead of the 2020 U.S. presidential election Five Thirty Eight (2020), from the perspective of the Republican party. (For the Democrats, skewness would have the opposite sign and smoothness would be the same.) Higher margins are more likely in states with more rugged and left-skewed polling distributions. Our model implies that parties should pursue more ambitious (albeit less risky, given the left-skew) political strategies in such states.
For example, pre-election polling distributions in 2020 would have suggested that Republicans had a higher chance of outperforming their mean polling margins in Minnesota and South Carolina than they did in neighboring Wisconsin and Georgia (Fig. 5c). This suggests that the Republicans had a greater incentive to pursue ambitious campaign strategies in these states, all else equal. Of course, there are many other factors, besides expected achievable vote margins, that are relevant in allocating campaign resources to states in practice. Such factors include (but are not limited to) the Electoral College rewards and how close the expected vote margins are to zero, given the first-past-the-post system in most states.
@@
Figure 5: Panel a: U.S. growth rates in gross domestic product (GDP) per capita (gray) Our World in Data (2024), compared moving means (black) and medians (dashed). The medians are greater than the means, indicating left skew. Panel b: distributions of private wealth among the 2500 top companies Dhiman (2024) and billionaires Elgiriyewithana (2023). The means are much greater than the medians, indicating right skew. Panel c: monthly polling margins, from the perspective of Republicans (i.e. Republican share -- Democrat share) in U.S. swing states during 2020, prior to the presidential election Five Thirty Eight (2020). Darker colors indicate larger average differences between the two parties. Lighter-colored states have smaller average margins.
@@ III Discussion
Our model formalizes the folk intuition that optimal ambition is both finite and aims to do better than average (Fig. 2a). We show empirical examples (from dating and college applications) in which people seem to behave as though they understand this, either consciously or sub-consciously (Fig. 2b,c). Although our model assumes that all agents face the same reward distribution, the available rewards—and consequently, the optimal ambition—varies widely across people and contexts in reality.
@@
Our model then adds more precise insights. First, we show that being too unsatisfiable is costlier than being too easily satisfied, on average (Fig. 2a). Second, we show that left-skewed and rugged reward distributions should motivate more ambition (Fig. 3). This exposes a subtle difference between optimal ambition and optimal risk taking, as left-skewed rewards motivate less risk-taking Benuzzi and Ploner (2024). Third, we show that upward social comparison is costly by creating unrealistic expectations (Fig. 4). We describe how these insights apply to real-world contexts relating to economic policy, wealth building, and political campaigns (Fig. 5). We discuss additional real-world examples in the introduction.
@@
Our model is admittedly simple; indeed, its accessibility to a broad audience is one of its contributions. However, ‘explore-exploit’ and sequential searches are common abstractions in areas such as microeconomics Kohn and Shavell (1974), management science Lavie et al. (2010); Greve (2007), harvesting Richard et al. (2018), and animal foraging MacArthur and Pianka (1966); Cowie (1977). Our ‘satisfaction’ framework is somewhat related to the satisficing concept in behavioral economics Simon (1955), but it is also distinct in that our threshold is optimized, rather than being a behavioral heuristic.
The framework of optimal search has been previously studied through different lenses in different disciplines, some of which provide insights into how our results might extend to more complex settings. For example, Kohn and Shavell Kohn and Shavell (1974) consider how time preference, risk aversion, uncertainty, or search costs affect optimal search. Their model is too complex for optimal satisfaction thresholds to be derived in general, but they show that risk aversion, search costs, and discounting future rewards would each reduce optimal satisfaction thresholds. These insights are intuitive. They also show that uncertainty can raise optimal satisfaction thresholds by increasing the information value of searching Kohn and Shavell (1974). This finding is somewhat counterintuitive in light of our finding that higher-than-optimal thresholds are costlier than lower-than-optimal thresholds (Fig. 2a), combined with that fact that left-skewed uncertainty typically motivates risk avoidance Benuzzi and Ploner (2024).
Applied mathematicians have widely studied what they call ‘the secretary problem’, where one seeks to maximize the chance of selecting the best option from a sequentially observed set Freeman (1983). For example, MacQueen and Miller MacQueen and Miller Jr (1960) characterize the optimal search duration under these conditions. The exact bounds and optimal strategies depend on the underlying distribution and corresponding assumptions. In the case of AR1 processes (like our model), several bounds have been derived for exponentially distributed steps  Novikov (2009); Christensen (2012); Christensen et al. (2011). However, because of their complexities and sensitivities to assumptions, empirical applications often rely on powerful but opaque machine learning approaches Becker et al. (2019); Ciocan and Mišić (2022).
Social comparisons have been theoretically predicted and empirically documented to affect decision-making in numerous human and animal contexts. Humans evaluate their successes and status relative to their peers Layard (2006, 2011), and, as a result, they make consumption decisions aimed at optimizing both absolute and relative outcomes (e.g., ref. Arrow and Dasgupta (2009)). Human fact-finding and decision-making processes can be either enhanced (e.g., the wisdom of crowds Surowiecki (2005)) or corrupted (e.g., groupthink Janis (1971)) by social forces. In animals, instincts for social conformism have also been linked to improved (e.g., ref. Berdahl et al. (2013)) and diminished (e.g., ref. Dussutour et al. (2007)) decision-making in different contexts.
@@
Future research should explore other complexities that our model does not consider. For example, there can be inherent benefits in reality to sticking with one strategy (e.g., accruing expertise, deepening a relationship). Similarly to search costs (e.g., disruptions from moving cities or changing jobs), this would likely reduce optimal satisfaction thresholds. Real people are also prone to weighing potential losses more than potential gains (known as ‘loss aversion’ Kahneman et al. (1991)). This, too, would reduce the optimal satisfaction threshold in absolute terms in our model, by reducing the utility that agents assign to low-end nominal rewards. Our agents are simplified in terms of both their decisions and their inferences about available rewards. Our social comparison model does not consider structured social networks. We model random rather than intelligent search processes, though it is straightforward to conjecture that intelligent search would increase optimal ambition by making high but rare rewards easier to find. There are many other worthwhile extensions to our framework.
@@
Despite these limitations, our conceptual results are intuitive and likely to generalize. For example, increasing ambition makes satisfactory rewards harder to find under any search process. Making high rewards more abundant—via skewness or ruggedness—makes ambitious targets easier to achieve. Upward social comparison distorts perceptions of reality, which hinders rational decision making.
@@
How ambitious to be is an important question in a wide range of professional, political, and personal contexts. Our model, results, and examples offer entry points for mathematically precise, but also accessible and intuitive, explorations of this question.
IV Methods